Circular No. 6010

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

PERIODIC COMET SHOEMAKER-LEVY 9 (1993e) For the Hubble Space Telescope (HST) observing team, K. S. Noll and

T. E. Smith, Space Telescope Science Institute, report: “HST observations on May 17.65 UT confirm the presence of an object between nuclei G (= 15) and F (= 16), as reported previously by Jewitt and Trentham (IAUC 5999). The HST WF image clearly shows an extended object located 4′′.20 east and 4′′.18 north of nucleus G. On May 17, the peak pixel of the new object is 39 times fainter than the peak pixel of nucleus G. The ratio of summed pixels in a 0′′.5 × 0′′.5 box, centered on nucleus G and the fainter new object, is 25 (3.5 magnitudes). The magnitude of nucleus G in this same aperture is R = 21.62 ± 0.05. HST PC images taken on Mar. 29.00 also show a faint object northeast of nucleus G. The object is extended and appears in each of the four exposures. The object is centered 1′′.85 east and 3′′.63 north of nucleus G. The ratio of the peak pixel in nucleus G to the peak pixel in this object is 24 and the ratio in a 0′′.5 × 0′′.5 box is 27. The brightness of nucleus G was constant between the Mar. 29 and May 17 images to within a few percent. The search for other objects near the limit of detectability has not been completed.”

1986 DA M. J. S. Belton, National Optical Astronomy Observatories, writes:

“1986 DA was determined to be an M-type near-earth minor planet and was studied in detail by radar and optical techniques during its only apparition. Thought to be a fragment of a differentiated asteroidal core, 1986 DA is easily reachable by spacecraft with small launch vehicles; it is therefore an important target for certain NASA Discovery missions that are currently being planned. However, the orbit is somewhat uncertain and requires new observations for orbit improvement to a level that would make it an acceptable target. 1986 DA will be at opposition in late Nov. 1994 at mag 20, passing through a phase angle of only 0^{o}.14 on Nov. 23.6 UT. I should appreciate receiving early notification of any observations that are made at this opposition, since this would strengthen the case for 1986 DA as a target for a space mission. An ephemeris can be found in the Catalogue of High-Precision Orbits of Unnumbered Minor Planets 1994 (see MPC 23363).”

1994 June 24 Daniel W. E. Green